Required information

[The following information applies to the questions displayed below.]

The following information is departmental cost allocation with two service departments and two production departments.

Percentage Service Provided to

Department Cost S1 S2 P1 P2
Service 1 (S1) $ 44,000 0 % 20 % 40 % 40 %
Service 2 (S2) 34,000 20 0 20 60
Production 1 (P1) 240,000
Production 2 (P2) 290,000

What is the amount of service department cost allocated to P1 and P2 using the direct method if the cost in P1 is changed from $240,000 to $260,000?

Total Cost Allocated
P1
P2
Answer #1

Answer : Total Cost Allocated to P1 is $30500 and to P2 is $47500

Total cost of S1 to be allocated is 44000. S1 provides 40% service to P1 and 40% to P2. Hence cost of S1 allocated to P1 and P2 is as below

Allocation to P1 = (40)/(40+40)*44000 = $22000

Allocation to P2 = (40)/(40+40)*44000 = $22000

Total cost of S2 to be allocated is 34000. S1 provides 20% service to P1 and 60% to P2. Hence cost of S2 allocated to P1 and P2 is as below

Allocation to P1 = (20)/(20+60)*34000 = $8500

Allocation to P2 = (60)/(20+60)*34000 = $25500

Hence total cost allocation to P1 = 22000+8500 = $30500

Hence total cost allocation to P2 = 22000+25500 = $47500